OpenStudy (anonymous):

What is the solution of the equation x − 8 = 17? A. -9 B. 9 C. 25 D. 27

OpenStudy (ivettef365):

Just add 8 to both sides

OpenStudy (anonymous):

x=25

OpenStudy (ivettef365):

x -8 = 17 +8 +8 -----------

OpenStudy (ivettef365):

Good job

OpenStudy (anonymous):

C)25

OpenStudy (anonymous):

Flip the equation around, so, 17+8=25. x = 25